# The Power of Iterative Filtering for Supervised Learning with (Heavy) Contamination

### v1

@@ 3 Iterative Polynomial Filtering
Let $\mathcal{D}=\mathcal{D}^{*}$. Fix any polynomial $p$ with $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[|p(\mathbf{x})|]\leq\epsilon/(4R)$. Due to the hypercontractivity of $\mathcal{D}$ (Definition 3.1), we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[(p(\mathbf{x}))^{2}]\leq(2A)^{2\ell}=\beta/2$, since $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[|p(\mathbf{x})|]\leq 1$. By Lemma A.1, since $m_{\mathrm{ref}}$ is large enough, w.p. at least $1-\frac{\delta}{2}$ over $S_{\mathrm{ref}}$, we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[(p(\mathbf{x}))^{2}]\leq\beta$.
It remains to show that with probability at least $21-\delta/$ 2, we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[|p(\mathbf{x})|]\leq\gamma=\epsilon/(2R)$. The following bound can be obtained for any $t\geq 1$ by applying the Marcinkiewicz–Zygmund inequality (see [43]), due to the fact that $S_{\mathrm{ref}}$ consists of $m_{\mathrm{ref}}$ i.i.d. examples.
@@
This concludes the proof of Theorem 3.2. ∎
@@ 4.1 Learning with Bounded Contamination
Our main result additionally requires that the target marginal distribution is hypercontractive. This assumption is inherited by Theorem 3.2.
@@
In Table 2, we summarize some of the new results we obtain as corollaries of Theorem 4.2, combined with appropriate known bounds on the approximate degree (see Section C.1). We also present the previous state-of-the-art results for comparison. For monotone functions and convex sets, no non-trivial results were known before this work. For halfspace intersections, we obtain exponential improvements over prior work, and for polynomial threshold functions, we obtain the first near-optimal error bounds. For small-depth circuits, we obtain an improved dependence on the depth.
@@
The proof of Theorem 4.2 is based on a delicate analysis of the error on the filtered dataset obtained by applying Theorem 3.2. The filtering algorithm is used to preserve the following important property of any polynomial $p^{*}$ that approximates a boolean function:
@@
We show that obtaining a filtered set that preserves this property is sufficient for learning with bounded contamination. To prove this, we crucially use the fact that the noise is bounded and that the right hand side of Eq. (4.1) is approximately equal to $1$ (rather than some larger constant). On the other hand, by applying part 2 of Theorem 3.2 on the polynomial $q(\mathbf{x})=(p^{*}(\mathbf{x}))^{2}-1$, we are able to ensure the desired property for $p^{*}$ on the filtered set. We give the full proof of our main result on learning with bounded contamination (Theorem 4.2) below.
Proof of Theorem 4.2.
@@
    First, the algorithm runs the filtering procedure of Theorem 3.2 (that is, Algorithm 1) on input $(S_{\mathrm{inp}},S_{\mathrm{ref}},m=|S_{\mathrm{inp}}|,2\ell,R=2,24\epsilon^{2}/{\sqrt{C}})$ to form the filtered dataset $\bar{S}_{\mathrm{filt}}$, where the labels are consistent with $\bar{S}_{\mathrm{inp}}$.
@@
    The algorithm outputs $h(\mathbf{x})=\operatorname{sign}(\widehat{p}(\mathbf{x})+\widehat{\tau})$, where $\widehat{\tau}\in{\mathbb{R}}$ minimizes the one-dimensional objective $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}_{\mathrm{filt}}}[y\neq\operatorname{sign}(\widehat{p}(\mathbf{x})+\tau)]$ over $\tau\in{\mathbb{R}}$.
We will now bound the error of $h$ on the clean dataset $\bar{S}_{\mathrm{cln}}$ according to which $\bar{S}_{\mathrm{inp}}$ was formed (see Definition 1.1). This suffices because, due to standard VC theory, and since $h$ is a polynomial threshold function of degree at most $\ell$, as long as $m=|\bar{S}_{\mathrm{cln}}|\geq C^{\prime}\frac{(d+1)^{\ell}+\log\frac{1}{\delta}}{\epsilon^{2}}$ for some sufficiently large universal constant $C^{\prime}\geq 1$, we have that $\operatorname*{\mathbb{P}}_{\mathbf{x}\sim\mathcal{D}}[f^{*}(\mathbf{x})\neq h(\mathbf{x})]$ is approximately equal to its empirical counterpart $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}_{\mathrm{inp}}}[y\neq h(\mathbf{x})]$.
Consider $p^{*}$ to be a polynomial of degree at most $\ell$ such that $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[(f^{*}(\mathbf{x})-p^{*}(\mathbf{x}))^{2}]\leq\epsilon^{\prime}$, where $\epsilon^{\prime}=\epsilon^{4}/{C}$. Since $p^{*}$ approximates a Boolean-valued function $f^{*}$, the values of $p^{*}$ should, in expectation, be close to either $1$ or $-1$. In particular, we obtain the following bound using the Cauchy-Schwarz inequality:
@@
Note that $q(\mathbf{x})=(p^{*}(\mathbf{x}))^{2}-1$ is a polynomial of degree at most $2\ell$. Therefore, Algorithm 1 can be used to filter out the points $\mathbf{x}$ such that $q(\mathbf{x})$ is too large and, in particular, due to Theorem 3.2, the following holds with probability at least $41-\delta/$ 4 over the random choice of $S_{\mathrm{ref}}$ and $\bar{S}_{\mathrm{cln}}$ as long as $m\geq C^{\prime}\frac{(2A(d+1))^{4\ell}}{\epsilon^{6}}\log\frac{1}{\delta}$ and $m_{\mathrm{ref}}\geq\frac{(C^{\prime}Ad)^{4\ell}}{\epsilon^{6}}(\log\frac{1}{\delta})^{8\ell+1}$:
@@
For the first term, we choose $m=|\bar{S}_{\mathrm{inp}}|=|\bar{S}_{\mathrm{cln}}|$ and we have $\frac{1}{2m}\sum_{(\mathbf{x},y)\in\bar{S}_{\mathrm{cln}}}|f^{*}(\mathbf{x})-p^{*}(\mathbf{x})|=\frac{1}{2}\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{cln}}}[|f^{*}(\mathbf{x})-p^{*}(\mathbf{x})|]$. In order to bound this quantity, we use the fact that $S_{\mathrm{cln}}$ is a set of $m$ i.i.d. samples from a hypercontractive distribution, combining the Marcinkiewicz-Zygmund inequality (which is a generalization of Chebyshev’s inequality to higher moments), hypercontractivity and boundedness of $f^{*}$ to obtain that, as long as $m\geq\frac{1}{\epsilon^{2}}(C^{\prime}A)^{2\ell}(\log(1/\delta))^{2\ell+1}$, with probability at least $41-\delta/$ 4 over the choice of $S_{\mathrm{cln}}$ we have:
@@
(4.3)${{\leq}}\frac{\epsilon^{2}m}{C^{\prime\prime}|S_{\mathrm{filt}}\setminus S_{\mathrm{cln}}|}+\frac{1}{|S_{\mathrm{filt}}\setminus S_{\mathrm{cln}}|}\sum_{\mathbf{x}\in S_{\mathrm{filt}}\cap S_{\mathrm{cln}}}|q(\mathbf{x})|$
@@
The first inequality in the expression above follows from the guarantee of Eq. (4.3), which is provided by the filtering algorithm. We may now use the Marcinkiewicz-Zygmund inequality once more to obtain that, as long as $m\geq\frac{1}{\epsilon^{4}}(C^{\prime}A)^{4\ell}(\log(1/\delta))^{4\ell+1}$, with probability at least $41-\delta/$ 4, we have:
@@
with probability at least $41-3\delta/$ 4 over the random choice of $\bar{S}_{\mathrm{inp}}$ and $S_{\mathrm{ref}}$.
We will now bound $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}_{\mathrm{inp}}}[y\neq h(\mathbf{x})]$ by $62\eta+5\epsilon/$ 6. We have the following:
@@
In order to complete the proof, we use the guarantee of Algorithm 1 that with probability at least $41-\delta/$ 4, we have
@@
Therefore, we have $|(\bar{S}_{\mathrm{cln}}\cap\bar{S}_{\mathrm{inp}})\setminus\bar{S}_{\mathrm{filt}}|\leq|(\bar{S}_{\mathrm{inp}}\setminus\bar{S}_{\mathrm{cln}})\setminus\bar{S}_{\mathrm{filt}}|+\epsilon m/6$. Note that the union of the sets $(\bar{S}_{\mathrm{inp}}\setminus\bar{S}_{\mathrm{cln}})\setminus\bar{S}_{\mathrm{filt}}$ and $\bar{S}_{\mathrm{filt}}\setminus\bar{S}_{\mathrm{cln}}$ is disjoint and equals to the set of adversarial examples $\bar{S}_{\mathrm{adv}}=\bar{S}_{\mathrm{inp}}\setminus\bar{S}_{\mathrm{cln}}$. By the definition of the noise model (Definition 1.1), we have that $|\bar{S}_{\mathrm{adv}}|\leq\eta m$ and $|\bar{S}_{\mathrm{cln}}\setminus\bar{S}_{\mathrm{inp}}|=|\bar{S}_{\mathrm{adv}}|\leq\eta m$. Therefore, we overall obtain the desired bound of
@@
This concludes the proof of Theorem 4.2. ∎
@@ 4.2 Learning with Heavy Contamination
In Table 3, we provide a number of end-to-end results for learning with heavy contamination, which are obtained by combining Theorem 4.4 with appropriate known bounds on the sandwiching degree (see Section C.2). The first row of Table 3 gives the bound $d^{\tilde{O}(Q^{2}k^{6}/\epsilon^{4})}$ for learning intersections of $k$ halfspaces with $Q$-heavy contamination by choosing $s=t=k$, as well as the bound $d^{\tilde{O}(Q^{2}k^{4}4^{k}/\epsilon^{4})}$ for arbitrary functions of $k$ halfspaces, with the choice $s=2^{k}$, $t=k$.
@@
Degree-$2$ PTFs $\operatorname{Unif}_{d}$ $d^{\tilde{O}(Q^{9}/\epsilon^{18})}$
@@
The proof of Theorem 4.4 uses once more Algorithm 1 to filter the input set, but this time we set the hyperparameter $R$ to a value that scales with the heavy contamination ratio $Q$. This ensures that we keep most of the clean points, because the filtering process is highly selective. Due to the choice of $R$, the required sandwiching degree scales with $Q$ as well. As a consequence, for many classes, our bounds are exponential in $Q$. After filtering, the approximation property of the sandwiching polynomials is preserved under the empirical distribution over the filtered set, and this is sufficient for efficient learnability of the optimum hypothesis on the filtered set, due to Theorem A.3 by [68]. We give the full proof of our main result on learning with heavy contamination (Theorem 4.4) below.
Proof of Theorem 4.4.
@@
    First, the algorithm runs the filtering procedure of Theorem 3.2 (that is, Algorithm 1) on input $(S_{\mathrm{inp}},S_{\mathrm{ref}},m=\frac{Q^{2}(C^{\prime}Ad)^{2\ell}\log\frac{1}{\delta}}{\epsilon^{5}},\ell,R=\frac{2Q}{\epsilon},\epsilon/3)$ to form the filtered dataset $\bar{S}_{\mathrm{filt}}$.
@@
    The algorithm outputs $h(\mathbf{x})=\operatorname{sign}(\widehat{p}(\mathbf{x})+\widehat{\tau})$, where $\widehat{\tau}\in{\mathbb{R}}$ minimizes the one-dimensional objective $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}_{\mathrm{filt}}}[y\neq\operatorname{sign}(\widehat{p}(\mathbf{x})+\tau)]$ over $\tau\in{\mathbb{R}}$.
@@
The second term in the bound above can be bounded by $\epsilon/3$, due to part 1 of Theorem 3.2. It remains to bound the first term.
Consider $f^{*}\in\mathcal{C}$ to be the concept in class $\mathcal{C}$ with minimum error on $\bar{S}_{\mathrm{filt}}$, i.e., we have that $f^{*}=\arg\min_{f\in\mathcal{C}}\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\in\bar{S}_{\mathrm{filt}}}[y\neq f(\mathbf{x})]$. Note that there is some $f^{\prime}\in\mathcal{C}$ that makes at most $M\cdot\mathsf{opt}_{\mathrm{total}}$ errors on the whole input dataset $\bar{S}_{\mathrm{inp}}$, due to the definition of $\mathsf{opt}_{\mathrm{total}}$ (see Definition 1.4). We have that $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\in\bar{S}_{\mathrm{filt}}}[y\neq f^{*}(\mathbf{x})]\leq\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\in\bar{S}_{\mathrm{filt}}}[y\neq f^{\prime}(\mathbf{x})]\leq{M}\cdot\mathsf{opt}_{\mathrm{total}}/{|\bar{S}_{\mathrm{filt}}|}$.
Moreover, let $p_{\mathrm{up}},p_{\mathrm{down}}$ be the $\frac{\epsilon^{2}}{72Q}$-sandwiching polynomials for $f^{*}$ and $p=p_{\mathrm{up}}-p_{\mathrm{down}}$, which is a non-negative polynomial with $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[|p(\mathbf{x})|]\leq\frac{\epsilon^{2}}{72Q}$. Due to part 2 of Theorem 3.2, the expectation of $p$ under $S_{\mathrm{filt}}$ is at most $\frac{\epsilon m}{3|\bar{S}_{\mathrm{filt}}|}$, i.e., $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{filt}}}[p_{\mathrm{up}}(\mathbf{x})-p_{\mathrm{down}}(\mathbf{x})]\leq\frac{\epsilon m}{3|\bar{S}_{\mathrm{filt}}|}$. Since $p_{\mathrm{up}}\geq f^{*}\geq p_{\mathrm{down}}$, we also have that
@@
We now observe that $h$ is the output of the $\mathcal{L}_{1}$ polynomial regression algorithm on $\bar{S}_{\mathrm{filt}}$ and, therefore, according to Theorem A.3 by [68], we overall have the following bound
@@
  • Learning with heavy contamination can be thought of as a version of testable agnostic learning (Definition 4.7), where the algorithm is asked to find a subset of the input that is structured enough, instead of merely verifying whether the whole input is structured.

### v2

@@ 3 Iterative Polynomial Filtering
Let $\mathcal{D}=\mathcal{D}^{*}$. Fix any polynomial $p$ with $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[|p(\mathbf{x})|]\leq\epsilon/(4R)$. Due to the hypercontractivity of $\mathcal{D}$ (Definition˜3.1), we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[(p(\mathbf{x}))^{2}]\leq(2A)^{2\ell}=\beta/2$, since $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[|p(\mathbf{x})|]\leq 1$. By Lemma˜A.1, since $m_{\mathrm{ref}}$ is large enough, w.p. at least $1-\frac{\delta}{2}$ over $S_{\mathrm{ref}}$, we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[(p(\mathbf{x}))^{2}]\leq\beta$.
It remains to show that with probability at least $1-\delta/2$, we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[|p(\mathbf{x})|]\leq\gamma=\epsilon/(2R)$. The following bound can be obtained for any $t\geq 1$ by applying the Marcinkiewicz–Zygmund inequality (see [46]), due to the fact that $S_{\mathrm{ref}}$ consists of $m_{\mathrm{ref}}$ i.i.d. examples.
@@
This concludes the proof of Theorem˜3.2. ∎
@@ 4.1 Learning with Bounded Contamination
Our main result additionally requires that the target marginal distribution is hypercontractive. This assumption is inherited by Theorem˜3.2.
@@
In Table˜2, we summarize some of the new results we obtain as corollaries of Theorem˜4.2, combined with appropriate known bounds on the approximate degree (see Section˜C.1). We also present the previous state-of-the-art results for comparison. For monotone functions and convex sets, no non-trivial results were known before this work. For halfspace intersections, we obtain exponential improvements over prior work, and for polynomial threshold functions, we obtain the first near-optimal error bounds. For small-depth circuits, we obtain an improved dependence on the depth.
@@
The proof of Theorem˜4.2 is based on a delicate analysis of the error on the filtered dataset obtained by applying Theorem˜3.2. The filtering algorithm is used to preserve the following important property of any polynomial $p^{*}$ that approximates a boolean function:
@@
We show that obtaining a filtered set that preserves this property is sufficient for learning with bounded contamination. To prove this, we crucially use the fact that the noise is bounded and that the right hand side of Eq. (4.1) is approximately equal to $1$ (rather than some larger constant). On the other hand, by applying part 2 of Theorem˜3.2 on the polynomial $q(\mathbf{x})=(p^{*}(\mathbf{x}))^{2}-1$, we are able to ensure the desired property for $p^{*}$ on the filtered set. We give the full proof of our main result on learning with bounded contamination (Theorem˜4.2) below.
Proof of Theorem˜4.2.
@@
    First, the algorithm runs the filtering procedure of Theorem˜3.2 (that is, Algorithm˜1) on input $(S_{\mathrm{inp}},S_{\mathrm{ref}},m=|S_{\mathrm{inp}}|,2\ell,R=2,24\epsilon^{2}/{\sqrt{C}})$ to form the filtered dataset $\bar{S}_{\mathrm{filt}}$, where the labels are consistent with $\bar{S}_{\mathrm{inp}}$.
@@
[…]
We will now bound the error of $h$ on the clean dataset $\bar{S}_{\mathrm{cln}}$ according to which $\bar{S}_{\mathrm{inp}}$ was formed (see Definition˜1.1). This suffices because, due to standard VC theory, and since $h$ is a polynomial threshold function of degree at most $\ell$, as long as $m=|\bar{S}_{\mathrm{cln}}|\geq C^{\prime}\frac{(d+1)^{\ell}+\log\frac{1}{\delta}}{\epsilon^{2}}$ for some sufficiently large universal constant $C^{\prime}\geq 1$, we have that $\operatorname*{\mathbb{P}}_{\mathbf{x}\sim\mathcal{D}}[f^{*}(\mathbf{x})\neq h(\mathbf{x})]$ is approximately equal to its empirical counterpart $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}_{\mathrm{inp}}}[y\neq h(\mathbf{x})]$.
Consider $p^{*}$ to be a polynomial of degree at most $\ell$ such that $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[(f^{*}(\mathbf{x})-p^{*}(\mathbf{x}))^{2}]\leq\epsilon^{\prime}$, where $\epsilon^{\prime}=\epsilon^{4}/{C}$. Since $p^{*}$ approximates a Boolean-valued function $f^{*}$, the values of $p^{*}$ should, in expectation, be close to either $1$ or 1-1. In particular, we obtain the following bound using the Cauchy-Schwarz inequality:
@@
Note that $q(\mathbf{x})=(p^{*}(\mathbf{x}))^{2}-1$ is a polynomial of degree at most $2\ell$. Therefore, Algorithm˜1 can be used to filter out the points $\mathbf{x}$ such that $q(\mathbf{x})$ is too large and, in particular, due to Theorem˜3.2, the following holds with probability at least $1-\delta/4$ over the random choice of $S_{\mathrm{ref}}$ and $\bar{S}_{\mathrm{cln}}$ as long as $m\geq C^{\prime}\frac{(2A(d+1))^{4\ell}}{\epsilon^{6}}\log\frac{1}{\delta}$ and $m_{\mathrm{ref}}\geq\frac{(C^{\prime}Ad)^{4\ell}}{\epsilon^{6}}(\log\frac{1}{\delta})^{8\ell+1}$:
@@
For the first term, we choose $m=|\bar{S}_{\mathrm{inp}}|=|\bar{S}_{\mathrm{cln}}|$ and we have $\frac{1}{2m}\sum_{(\mathbf{x},y)\in\bar{S}_{\mathrm{cln}}}|f^{*}(\mathbf{x})-p^{*}(\mathbf{x})|=\frac{1}{2}\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{cln}}}[|f^{*}(\mathbf{x})-p^{*}(\mathbf{x})|]$. In order to bound this quantity, we use the fact that $S_{\mathrm{cln}}$ is a set of $m$ i.i.d. samples from a hypercontractive distribution, combining the Marcinkiewicz-Zygmund inequality (which is a generalization of Chebyshev’s inequality to higher moments), hypercontractivity and boundedness of $f^{*}$ to obtain that, as long as $m\geq\frac{1}{\epsilon^{2}}(C^{\prime}A)^{2\ell}(\log(1/\delta))^{2\ell+1}$, with probability at least $1-\delta/4$ over the choice of $S_{\mathrm{cln}}$ we have:
@@
[…]
@@
The first inequality in the expression above follows from the guarantee of Eq. (4.3), which is provided by the filtering algorithm. We may now use the Marcinkiewicz-Zygmund inequality once more to obtain that, as long as $m\geq\frac{1}{\epsilon^{4}}(C^{\prime}A)^{4\ell}(\log(1/\delta))^{4\ell+1}$, with probability at least $1-\delta/4$, we have:
@@
with probability at least $1-3\delta/4$ over the random choice of $\bar{S}_{\mathrm{inp}}$ and $S_{\mathrm{ref}}$.
We will now bound $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}_{\mathrm{inp}}}[y\neq h(\mathbf{x})]$ by $2\eta+5\epsilon/6$. We have the following:
@@
In order to complete the proof, we use the guarantee of Algorithm˜1 that with probability at least $1-\delta/4$, we have
@@
Therefore, we have $|(\bar{S}_{\mathrm{cln}}\cap\bar{S}_{\mathrm{inp}})\setminus\bar{S}_{\mathrm{filt}}|\leq|(\bar{S}_{\mathrm{inp}}\setminus\bar{S}_{\mathrm{cln}})\setminus\bar{S}_{\mathrm{filt}}|+\epsilon m/6$. Note that the union of the sets $(\bar{S}_{\mathrm{inp}}\setminus\bar{S}_{\mathrm{cln}})\setminus\bar{S}_{\mathrm{filt}}$ and $\bar{S}_{\mathrm{filt}}\setminus\bar{S}_{\mathrm{cln}}$ is disjoint and equals to the set of adversarial examples $\bar{S}_{\mathrm{adv}}=\bar{S}_{\mathrm{inp}}\setminus\bar{S}_{\mathrm{cln}}$. By the definition of the noise model (Definition˜1.1), we have that $|\bar{S}_{\mathrm{adv}}|\leq\eta m$ and $|\bar{S}_{\mathrm{cln}}\setminus\bar{S}_{\mathrm{inp}}|=|\bar{S}_{\mathrm{adv}}|\leq\eta m$. Therefore, we overall obtain the desired bound of
@@
This concludes the proof of Theorem˜4.2. ∎
@@ 4.2 Learning with Heavy Contamination
In Table˜3, we provide a number of end-to-end results for learning with heavy contamination, which are obtained by combining Theorem˜4.4 with appropriate known bounds on the sandwiching degree (see Section˜C.2). The first row of Table˜3 gives the bound $d^{\tilde{O}(Q^{2}k^{6}/\epsilon^{4})}$ for learning intersections of $k$ halfspaces with $Q$-heavy contamination by choosing $s=t=k$, as well as the bound $d^{\tilde{O}(Q^{2}k^{4}4^{k}/\epsilon^{4})}$ for arbitrary functions of $k$ halfspaces, with the choice $s=2^{k}$, $t=k$.
@@
Degree-22 PTFs $\operatorname{Unif}_{d}$ $d^{\tilde{O}(Q^{9}/\epsilon^{18})}$
@@
The proof of Theorem˜4.4 uses once more Algorithm˜1 to filter the input set, but this time we set the hyperparameter $R$ to a value that scales with the heavy contamination ratio $Q$. This ensures that we keep most of the clean points, because the filtering process is highly selective. Due to the choice of $R$, the required sandwiching degree scales with $Q$ as well. As a consequence, for many classes, our bounds are exponential in $Q$. After filtering, the approximation property of the sandwiching polynomials is preserved under the empirical distribution over the filtered set, and this is sufficient for efficient learnability of the optimum hypothesis on the filtered set, due to Theorem˜A.3 by [60]. We give the full proof of our main result on learning with heavy contamination (Theorem˜4.4) below.
Proof of Theorem˜4.4.
@@
    First, the algorithm runs the filtering procedure of Theorem˜3.2 (that is, Algorithm˜1) on input $(S_{\mathrm{inp}},S_{\mathrm{ref}},m=\frac{Q^{2}(C^{\prime}Ad)^{2\ell}\log\frac{1}{\delta}}{\epsilon^{5}},\ell,R=\frac{2Q}{\epsilon},\epsilon/3)$ to form the filtered dataset $\bar{S}_{\mathrm{filt}}$.
@@
[…]
@@
The second term in the bound above can be bounded by $\epsilon/3$, due to part 1 of Theorem˜3.2. It remains to bound the first term.
Consider $f^{*}\in\mathcal{C}$ to be the concept in class $\mathcal{C}$ with minimum error on $\bar{S}_{\mathrm{filt}}$, i.e., we have that $f^{*}=\arg\min_{f\in\mathcal{C}}\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\in\bar{S}_{\mathrm{filt}}}[y\neq f(\mathbf{x})]$. Note that there is some $f^{\prime}\in\mathcal{C}$ that makes at most $M\cdot\mathsf{opt}_{\mathrm{total}}$ errors on the whole input dataset $\bar{S}_{\mathrm{inp}}$, due to the definition of $\mathsf{opt}_{\mathrm{total}}$ (see Definition˜1.4). We have that $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\in\bar{S}_{\mathrm{filt}}}[y\neq f^{*}(\mathbf{x})]\leq\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\in\bar{S}_{\mathrm{filt}}}[y\neq f^{\prime}(\mathbf{x})]\leq{M}\cdot\mathsf{opt}_{\mathrm{total}}/{|\bar{S}_{\mathrm{filt}}|}$.
Moreover, let $p_{\mathrm{up}},p_{\mathrm{down}}$ be the $\frac{\epsilon^{2}}{72Q}$-sandwiching polynomials for $f^{*}$ and $p=p_{\mathrm{up}}-p_{\mathrm{down}}$, which is a non-negative polynomial with $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[|p(\mathbf{x})|]\leq\frac{\epsilon^{2}}{72Q}$. Due to part 2 of Theorem˜3.2, the expectation of $p$ under $S_{\mathrm{filt}}$ is at most $\frac{\epsilon m}{3|\bar{S}_{\mathrm{filt}}|}$, i.e., $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{filt}}}[p_{\mathrm{up}}(\mathbf{x})-p_{\mathrm{down}}(\mathbf{x})]\leq\frac{\epsilon m}{3|\bar{S}_{\mathrm{filt}}|}$. Since $p_{\mathrm{up}}\geq f^{*}\geq p_{\mathrm{down}}$, we also have that
@@
We now observe that $h$ is the output of the $\mathcal{L}_{1}$ polynomial regression algorithm on $\bar{S}_{\mathrm{filt}}$ and, therefore, according to Theorem˜A.3 by [60], we overall have the following bound
@@
  • Learning with heavy contamination can be thought of as a version of testable agnostic learning (Definition˜4.7), where the algorithm is asked to find a subset of the input that is structured enough, instead of merely verifying whether the whole input is structured.